OpenStudy (anonymous):

Find the product of (3r^2)(2r^2 + r - 7)

OpenStudy (jackellyn):

=3r^2(2r^2) + 3r^2(r) - 3r^2(7) Can you do it now? Remember, when multiplying variables the exponents add together. (example: x^2(x)=x^3 because 2+1=3).

OpenStudy (anonymous):

thx

OpenStudy (jackellyn):

no problem :)